### v1

@@ 5 CONCLUSION AND FUTURE DIRECTIONS
The measurements presented in this paper represent the successful determination of CHIME correlator settings for drone-based beam calibration and demonstrate the first use of a drone as a reference calibrator for a large cylindrical radio interferometer. The comparisons between the drone, holography, and solar measurements indicate that the FWHM and centroid position are consistent across all three data sets. Moreover, the evolution of these parameters as a function of declination, frequency, and feed index are shown to behave similarly in all three data sets, despite differences in measurement technique. This shows that drone beam measurements are possible, even at the relatively slow data cadence of CHIME, by using its pulsar gating mode. The analysis also demonstrates the complexities of measuring a cylindrical array with a drone and highlighted the areas for improvement, in particular a higher precision constraint on drone altitude is essential for this mapping and designing flight trajectories which can more easily be mapped to comparison data sets.
During the same maintenance window, our group conducted nine other drone beam mapping test flights which are prime targets for future analysis. We performed similar East-to-West scans at 45 and 55 zenith angle, North-to-South transits along the focal line of Cylinder C, and gridded flights over the center of the array. In addition, test flights were performed where the drone rotated through 180 of heading (and thus, polarization) at several locations above the array to test the polarization response of the feeds as a function of angle.
@@
Now that the correlator firmware and pulsar timing parameter files have been generated, we can return to CHIME to conduct additional drone flights targeting regions of interest within the beam, including two-dimensional polarized properties of the beam maps.

### v2

@@ 5 CONCLUSION AND FUTURE DIRECTIONS
The measurements presented in this paper represent the successful development of a correlator data acquisition mode for drone-based beam calibration and demonstrate the first use of a drone as a reference calibrator for a large cylindrical radio interferometer. The comparisons between the drone, holography, and solar measurements indicate that the FWHM and centroid position are consistent across all three data sets. Moreover, the evolution of these parameters as a function of declination, frequency, and feed index are shown to behave similarly in all three data sets, despite differences in measurement technique. This shows that drone beam measurements are possible, even at the relatively slow data cadence of CHIME, by using its pulsar gating mode. The analysis also demonstrates the complexities of measuring a cylindrical array with a drone and highlighted the areas for improvement, in particular a higher precision constraint on drone altitude is essential for this mapping and designing flight trajectories which can more easily be mapped to comparison data sets.
During the same maintenance window, our group conducted nine other drone beam mapping test flights which are prime targets for future analysis. We performed similar East-to-West scans at 45 and $5$ zenith angle, North-to-South transits along the focal line of Cylinder C, and gridded flights over the center of the array. In addition, test flights were performed where the drone rotated through 180 of heading (and thus, polarization) at several locations above the array to test the polarization response of the feeds as a function of angle.
@@
Now that the correlator firmware and pulsar timing parameter files have been generated, we can also return to CHIME to conduct additional drone flights targeting regions of interest within the beam, including two-dimensional polarized properties of the beam maps. Some limitations noted in this paper are important to address in these future flights. First, neglecting the transmitter beam is currently a large sytematic error for angles beyond the main beam, and will need to be incorporated and removed for future work. Along these lines, making measurements equivalent to the more standard Ludwig-III polarization definition are important for future work.